Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 in the present but Y = 100 in the future. Let MC = 0. Let reserves = 200. Consider the basic two-period model. What is consumption of the resource in the present?

48.32

55.23

62.44

70.12

none of the above
Answer #1

Answer option 5)

in present, Q1= 150-P+2*50

Q1= 250-P

In future, Q2= 150-P+2*100

Q2= 350-P

So at eqm extraction path

250-Q1 = (350-Q2)/(1+.1)

275-1.1Q1 = 350-Q2

Q2-1.1Q1 = 75

put, Q1+Q2= 200

Q2= 200-Q1

So, 200-Q1-1.1Q1 = 75

125= 2.1Q1

so, Q1*= 59.52